The Second International Symposium on Parallel Symbolic Computation (PASCO’97) presents original research on all aspects of high performance symbolic computation. High performance is meant as pertaining to the solution of large problems by means of parallel, distributed, multi-processor, or networked computers. Topics include: high performance parallel algebraic computation, high performance combinatorial and discrete methods, high preformance automatic theorem proving, design of high performance symbolic, functional, and constraint/logic languages and systems. PASCO’97 will be held under the sponsorship of ACM SIGSAM and SIGNUM at the Aston Wailea Resort, Maui, Hawaii, U.S.A. from July 20 – 22, 1997. PASCO’97 will be held in federation with ISSAC’97 the International Symposium on Symbolic and Algebraic Computation. (July 20–23 at the same location), and the exhibitions, systems demonstration area, and social program will be shared among the conferences. PASCO’97 registration permits participation in all ISSAC’97 activities For more details about the conference and about registration, please consult the enclosed registration leaflet or the web page

http://www.csc.ncsu.edu/pasco97/, or contact the local organizers.

Symposium Chair: Hoon Hong RISC-Linz Johannes Kepler University A-4040 Linz, Austria

Phone: +43 (7236) 3231-48
Fax: +43 (7236) 3231-30

 email: Hoon.Hong@risc.uni-linz.ac.at

Local Arrangements: Stan Steinberg University of New Mexico

 email: stanly@math.unm.edu

Publicity: Markus Hitz North Carolina State University

 email: mahitz@eos.ncsu.edu

Registrations: Maria Paola Bonacina University of Iowa

 email: bonacina@cs.uiowa.edu

Treasurer: Sam Dooley IBM T. J. Watson Research

 email: dooley@watson.ibm.com

Program Committee:

Gene Cooperman  Northeastern U.
Wayne Eberly  U. Calgary
Tetsuro Fujise  Mitsubishi Research
Kevin Hammond  U. of St. Andrews
Jieh Hsiang  National Taiwan U.
Jean-Marie Jacquet FUNDP Namur
Sverker Janson  Swedish Inst. Comp. Sci.
Erich Kaltofen (chair) N. C. State University
Herbert Kuchen  RWTH Aachen
Wolfgang Küchlin  U. Tübingen
Mohamed O. Rayes Texas Instruments
Jean-Louis Roch  LMC/IMAG Grenoble
Stan Steinberg  U. New Mexico

Preliminary Program

Sunday, July 20

Session P, Track 1 (2:00pm–3:00pm)

PASCO ’97 begins at 2pm. Efficient Parallel Solution of Sparse Systems of Linear Diophantine Equations Mark Giesbrecht, University of Manitoba Fast Rectangular Matrix Multiplications and Improving Parallel Matrix Computations Xiaohan Huang and Victor Y. Pan, City University of New York Towards an efficient implementation of a fast algorithm for multipoint polynomial evaluation and its parallel processing Hirokazu Murao, University of Tokyo and Tetsuro Fujise, Mitsubishi Research Institute

Session Q, Track 1 (3:30pm–4:10pm)

Fast parallel algebraic numbers computations Thierry Gautier, ETH Zurich and Jean-Louis Roch, LMC-IMAG Grenoble Processor-Efficient Parallel Matrix Inversion over Abstract Fields: Two Extensions Wayne Eberly, University of Calgary

Monday, July 21

Session A, Track 2 (8am–9:30am)

An evaluation of Penny: a system for fine grain implicit parallelism Johan Montelius and Seif Haridi, Swedish Institute of Computer Science Evaluating Parallel Logic Programming Systems on Scalable Multiprocessors Vítor Santos Costa, Universidade do Porto Ricardo Bianchini and Inès de Castro Dutra, Federal University of Rio de Janeiro Extended Dynamic Dependent And-parallelism in ACE Gopal Gupta and Enrico Pontelli, New Mexico State University Distributed Partial Evaluation Michael Sperber and Peter Thiemann, Universität Tübingen

Session B, Track 2 (9:45am–11:15am)

Experiments with subdivisions of search in distributed theorem proving Maria Paola Bonacina, University of Iowa Deriving Efficient Parallel Programs for Complex Recurrences Wei-Ngan Chin, Y.M. Teo, and S.H. Tan, National University of Singapore A Parallel Implementation of Cost-based Abductive Reasoning Shohei Kato, Hirohisa Seki, and Hidenori Itoh, Nagoya Institute of Technology Parallel Implementation through Object Graph Rewriting John Glauert, University of East Anglia, Norwich

Session invited lecture, Track 1&2 (11:30am–12:30pm)

Title to be announced Monica Lam, Stanford University

Session C, Track 2 (2pm–3pm)

Computation of replicable functions on Risa/Asir Masayuki Noro, Fujitsu Laboratories Ltd. and John McKay, Concordia University, Montréal The computational complexity of the Wu-Ritt type unmixed decomposition of radical ideals Agnes Szanto, Cornell University Mixed representation of polynomials oriented towards fast parallel shift Eugene Zima, Moscow State University

Session D, Track 2 (4pm–5pm)

Improvement of Message Communication in Concurrent Logic Language Kazuhiko Ohno, Masahiko Ikawa, Shin-ichiro Mori, Hiroshi Nakashima, and Shinji Tomita, Kyoto University A Proposal for Syntactic Data Integration for Math Protocols Olaf Bachmann and Hans Schönemann, Universität Kaiserslautern and Simon Gray, Kent State University An Overview of the Design of Distributed Oz Seif Haridi and Peter Van Roy, Swedish Institute of Computer Science and Gert Smolka, Universität des Saarlandes

Tuesday, July 22

Session E, Track 2 (8am–9:30am)

Tools for Parallel/Distributed Mathematical Computation Paul S. Wang, Kent State University CABAL: Polynomial and power series algebra on a parallel computer Arthur Norman, Trinity College, Cambridge and John Fitch, University of Bath CALYPSO: a Computer Algebra Library for Parallel Symbolic Computation Giovanni Cesari, ETH Zurich and Università degli Studi di Trieste Maple on a massively parallel, distributed memory machine Laurent Bernardin, ETH Zurich PASCO ’97 ends.